In this unit, we learned about measurements and dimensional analysis, which is referred to as conversion factors. Answer the questions below.

Note: each question may have more than one part, so be sure thoroughly answer each part.

You must provide a logical explanation with sufficient detail to demonstrate mastery of the chemistry concept related to the question in order to pass. Although there is not a length requirement, answers less than 250 words will likely not offer sufficient detail to earn a passing grade.

I will need this completed within the next couple of days. Please no plagiarism and anyone willing to help will receive a very good review!

***KEEP IN MIND ANSWERS NEED TO BE OVER 250 WORDS PER QUESTION

  1. Describe the units of measure most commonly used in Chemistry for length, mass and volume. Include in your description the name and abbreviation for the metric system base units.  Describe a situation in the chemistry classroom when you might measure length, mass and volume. You would not need to use the same situation for all three quantities.
  2. Explain why using dimensional analysis to solve problems in Chemistry is a useful strategy.
  3. Write and solve your own 2-step dimensional analysis problem that utilizes a metric system conversion and a metric-U.S. System conversion. Add detailed narratives explaining each step.
  4. Explain the process and rational for using the water displacement method to measure the volume of a solid object. Explain how you should find the density of an object using the water displacement method. Then, write and solve your own math problem to find the density of an object for which the volume is found using the water displacement method.
Answer #1

Your Answer:

1) Describe the units of measurements:

For length:  

1) Standard unit of length is "Meter" which is denoted by 'm'.

2) A meter is then divided into 100 equal parts. Each part is called as centimeter and it is denoted as "cm". Therefore, 1 meter = 100 centimeters.

3) Long distances are measured in Kilometer. It is abbreviated as "Km".

Therefore, 1 Kilometer = 1000 meters.

Examples:  

We use scales in school (or) college levels that are either 15cms or 30 cms long to draw lines.

We can use tapes to measure length of blackboard etc.,

For Mass:

1) Standard unit of mass is Kilogram which is denoted by "Kg".

2) If we again divide Kg in 1000 parts, then each part is called "Gram" which is denoted by "g".

Therefore 1 Kilogram = 1000 grams.

3) Another u nit is "Quintal". 1 Quintal = 100 Kgs.

Examples:  

1) In chemistry laboratory, we measure amorphous chemicals either in grams (or) milligrams (or) Kilograms.

2) Even in class room also, we use to measure chalk in grams.

For Volume:

1) Volume is related to the quantity present.

2) It is measured in Cubic Centimeters in metric (Cm3) and also liter and milliliter are the other units to measure volume.

Examples:  

1) In chemistry laboratory, we measure liquid chemicals like aqueous NaOH either in milliliters (or) liters (or) Cm3.

2) Almost, every student in classroom carrying 1 liter water bottle.

----------------------------------------------------------------------------------------------------------------------------------Explain why using dimensional analysis to solve problems in Chemistry is a useful strategy.

Your Answer:

1) Dimensional Analysis is a problem solving method that uses the fact that any number can be multliplied by one without changing its value.

2) We can convert any unit to another unit of same dimension.

3) Hence using dimensional analysis to solve problem in chemistry is useful strategy.

Examples:

1) Convert 20.0 mL to L.

Ans: We know that 1000 mL = 1 L. Therefore 1 mL = 1/1000 L.

So, for 20 mL = 20/1000 L = 0.020 L.

----------------------------------------------------------------------------------------------------------------------------------

Your Answer:

Water displacement method to find volume of solid object:

  • We have taken a solid object i.e., stone and one jar filled with water at 1000mL level.
  • Weigh the stone and if we weigh suppose 500g.
  • Put the stone in water, water level rises to 1200 mL.
  • So we can calculate volume of stone = New water level - Original water level

= 1200 - 1000 = 200 mL.

  • This is how we can find the volume of a solid.
  • Once we know mass and volume of the solid object, we can find the density.
    • Density (d) = mass/volume.
    • Density (d) = 500g / 200 mL = 2.5 g / mL.
  • Since 1 mL = 1 Cubic Centimeter. (cc)
  • Therefore Density (d) = 2.5 g / cc.

----------------------------------------------------- XXXXXXX ----------------------------------------------------------